# SARATR-X: A Foundation Model for Synthetic Aperture Radar Images Target Recognition

### v1

@@ Abstract
Synthetic aperture radar (SAR) is essential in actively acquiring information for Earth observation. SAR Automatic Target Recognition (ATR) focuses on detecting and classifying various target categories under different image conditions. The current deep learning-based SAR ATR methods are typically designed for specific datasets and applications. Various target characteristics, scene background information, and sensor parameters across ATR datasets challenge the generalization of those methods. This paper aims to achieve general SAR ATR based on a foundation model with Self-Supervised Learning (SSL). Our motivation is to break through the specific dataset and condition limitations and obtain universal perceptual capabilities across the target, scene, and sensor. A foundation model named SARATR-X is proposed with the following four aspects: pre-training dataset, model backbone, SSL, and evaluation task. First, we integrated 14 datasets with various target categories and imaging conditions as a pre-training dataset. Second, different model backbones were discussed to find the most suitable approaches for remote-sensing images. Third, we applied two-stage training and SAR gradient features to ensure the diversity and scalability of SARATR-X. Finally, SARATR-X has achieved competitive and superior performance on 5 datasets with 8 task settings, which shows that the foundation model can achieve universal SAR ATR. We believe it’s time to embrace fundamental models for SAR image interpretation in the era of increasing big data.
@@
Synthetic Aperture Radar (SAR), Target Recognition, Foundation Model, Self-Supervised Learning (SSL), Deep Learning, Masked Image Modeling (MIM)
1 Introduction
Based on electromagnetic scattering in microwave frequency bands, Synthetic Aperture Radar (SAR) [62, 22, 53] is essential in actively acquiring information for Earth observation with stable imaging capacity under various weather and lighting conditions. SAR Automatic Target Recognition (ATR) aims to automatically localize and classify interested objects in SAR images, e.g., detection and classification. It has been investigated extensively with various civilian and military applications [69, 36, 73, 57]. In the past decade, deep learning has reinvigorated SAR ATR with impressive breakthroughs [31, 37, 15]. However, as shown in Fig 1 and 2, due to the extensive acquisition conditions as well as the high costs of collection and annotation, there are a plethora of special datasets and algorithms designed for different applications of SAR ATR. For example, nearly a dozen new target datasets and dozens of attention modules have been developed for SAR target recognition in the last few years. However, the target characteristics, scene background information and sensor parameters across datasets challenge their generalization. In contrast to specialized algorithms for particular applications, we aim to investigate a general SAR target recognition method.
A foundation model [5] pre-trained in a task-agnostic manner (generally via self-supervision learning) on extensive data can be flexibly adapted to a wide range of downstream tasks under various conditions. Self-supervised learning (SSL) [45, 4] can mitigate label inefficiency by exploring supervision directly from the data itself, thereby reducing reliance on expensive expert labeling and efficiently scaling data and models. Hence, SSL has the capability to leverage extensive SAR unlabeled data for constructing foundational models, which can achieve universal recognition with a few labeled target samples in different categories and conditions. As shown in Table I, foundation models are thriving in remote sensing with superior performance in a diverse range of modalities and tasks. In particular, multi-modality foundation models have recently appeared, indicating this research topic’s great potential. But, there lacks a foundation model that can effectively accomplish various tasks for SAR ATR, such as FG-MAE [79] and SkySense [26] mainly focus on SAR scene classification and segmentation, and our previous SAR-JEPA [38] and MSFA [43] explored the object classification and detection tasks separately. It is still necessary to explore the development of an analogous ATR foundation model for SAR image interpretation, which lies at the intersection of SAR technologies and frontier AI technologies (in particular, the big models).
@@ 1 Introduction
Figure 1: Motivation and performance of SARATR-X. (a) SAR ATR has various imaging conditions, including targets, scenes, and sensors. However, due to high cost, its datasets are collected under specific settings and tasks. For example, the MSTAR dataset [1] is a ten-type vehicle target classification dataset in X-band and grass scenarios, while SAR-Aircraft is a seven-type aircraft target detection dataset in three airports and a C-band satellite. Many specialized algorithms have been proposed for these datasets. However, the different target characteristics, scene background information and sensor parameters between these datasets challenge the generalization of existing algorithms. This paper aims to explore a SAR ATR foundation model, a general method to achieve various tasks. (b) Our SARATR-X performs well on 5 datasets with 8 settings compared with previous methods. For target classification, SARATR-X is better than the previous SSL method (BIDFC [92]) on fine-grained vehicle MSTAR dataset [1] with the few-shot setting. Besides, it performs well under extended operating conditions (EOCs) [94], i.e., the imaging condition variation in depression angle (EOCs-Depression), target configuration (EOCs-Config), and version (EOCs-Version). For object detection, SARATR-X demonstrates competitive performance with previous carefully designed supervised methods on various categories (SARDet-100K [43] and OGSOD [71]) and specific categories on ship (SSDD [97]), and aircraft (SAR-AIRcraft [80]). Our study shows the potential of the foundation model for universal SAR target recognition.
Thanks to the development of SAR sensors and the enormous efforts of researchers, numerous SAR target datasets have emerged over the past five years in Fig. 2. Despite image annotation remaining a challenging and expensive endeavor, SAR sensors have produced vast amounts of data at a significantly accelerated pace compared to the previous ten years. Meanwhile, the investigation of SSL and foundation models for SAR ATR is advancing with increasing datasets. As shown in Table I, existing research [92, 79, 56, 38, 43] have discussed different factors of a foundation model, and these studies are progressively advancing towards a goal of achieving general SAR ATR. For example, BIDFC [92] performs much better than supervised methods in the few-shot classification of the MSTAR vehicle dataset, SAR-JEPA [38] demonstrates that SSL can effectively improve the classification accuracy for different targets across datasets, and MSFA [43] shows improved results of ship detection. Inspired by their insightful work, we formally propose a SAR ATR foundation model, SARATR-X, to achieve a breakthrough in universal SAR target recognition.
In this work, we lay the groundwork for developing SARATR-X, a foundation model specifically tailored for SAR ATR tasks. It aims to learn generalizable representations from unlabeled SAR images, offering a basis for efficient adaptation across various downstream ATR tasks. The four parts for creating SARATR-X include pre-training datasets, model backbones, self-supervised learning, and evaluation tasks.
@@
Pre-training datasets need to contain diverse target categories and imaging conditions to accommodate various downstream tasks. However, SAR ATR lacks a large-scale dataset like ImageNet [19], and the most used MSTAR dataset in Table II only includes fine-grained vehicle categories, which is not suitable for larger-scale pre-training. Therefore, many SSL methods combine different datasets; in particular, SARDet-100K incorporates [43] 9 SAR target detection datasets. With the increasing number of open-source datasets for SAR ATR, we integrate the vast majority of them for pre-training. As shown in Fig 2 and Table III, we apply 14 classification and detection datasets with different target categories and imaging conditions to explore the foundation model’s potential.
@@
Model backbones aim to achieve better spatial representation in remote sensing images, especially small target signatures in large imagery. As shown in Table I, researchers used two architectures (Transformer or convolutional neural network). Transformer has better spatial resolution without downsampling, and Convolutional Neural Network (CNN) has higher efficiency with its convolution kernel. MSFA discussed the effectiveness of different architectures, and the results showed that the Swin Transformer backbone performs better. Therefore, we discuss different model backbones to find the most suitable approaches for the properties of remote sensing images. HiViT [99] is our choice, which combines Swin Transformer advantages and can drop patches in MIM.
@@
Self-supervised learning faces the challenge of SAR image quality. SAR images contain speckle noise due to coherent imaging, and their visual features are not as distinct and rich as RGB images. As shown in Table II, contrastive learning [92, 56] uses data augmentation and pre-processing to reduce noise, while MIM [79, 38, 43] has a pixel-to-feature improvement for guide signals. Therefore, we consider that the problem to be solved by SAR SSL is how to construct high-quality guide signals. For example, PGIL [29] leveraged a sub-frequency feature of SAR complex images to learn physics information, and our SAR-JEPA [38] applied multi-scale gradient ratio to solve multiplicative speckle noise and capture target shape. Furthermore, multi-stage training [43] from ImageNet to SAR diminished noise interference on model diversity in Fig. 5. Therefore, we apply two-stage training from ImageNet to SAR and use multi-scale gradient features as high-quality guide signals for SAR MIM.
@@
Evaluation tasks need to comprehensively evaluate the performance of the foundation model under different tasks and settings. Benefiting from the open-source target datasets, we first construct a fine-grained classification dataset with various categories to evaluate the effectiveness of the proposed improvements. In the end, we perform a comprehensive comparison between the proposed SARATR-X and existing SOTA methods in public classification and detection tasks.
Our SARATR-X has achieved a competitive and superior performance on 5 datasets with 8 task settings in Fig 1. And its codes and weights can be found at https://github.com/waterdisappear/SARATR-X. We hope this work will advance the development of foundation models and general SAR target recognition. The main contributions are summarized as follows:
TABLE I: List of visual foundation models in remote sensing. There are many foundation models for various modalities and tasks, but research (FG-MAE, SkySense, and OFA-Net) focuses on scene-level tasks with the SAR modality. Based on our previous study SAR-JEPA and MSFA [38, 43], SARATR-X fills this gap in the foundation model for SAR ATR. These foundation models’ SSL methods are divided into two approaches: one is contrastive, such as contrastive learning, to obtain invariant features; another is generative, such as masked image modeling, to generate and predict features. SD: Stable Diffusion. Swin: Swin Transformer.
Foundation model Year Modality Dataset Backbone SSL Tasks
SatMAE [13] 2022 Multi-spectral, RGB fMoW RGB/Sentinel ViT Generative Scene classification and semantic segmentation
RVSA [77] 2022 RGB MillionAID ViT, ViTAE Generative Scene classification, object detection, and semantic segmentation
RingMo [65] 2022 RGB Self-built ViT, Swin Generative Scene classification, object detection, semantic segmentation, and change detectionn
RingMo-Sense [90] 2023 RGB Self-built Video Swin Generative object detection and tracking, images segmentation, and 3 prediction tasks
CMID [55] 2023 RGB MillionAID ResNet, Swin Generative & Contrastive Scene classification, object detection, and semantic segmentation
GFM [52] 2023 RGB GeoPile Swin Generative Scene classification, semantic segmentation, change detection, and super-resolution
DiffusionSat [32] 2023 RGB fMoW, Satlas, SpaceNet SD Generative Image generation, super-resolution, temporal generation, and in-painting
Scale-MAE [58] 2023 RGB fMoW RBG ViT Generative Scene classification, object detection, and semantic segmentation
FG-MAE [79] 2023 Multi-spectral, SAR SSL4EO-S12 ViT Generative Scene classification and semantic segmentation
SMLFR [17] 2024 RGB GeoSense ConvNeXt Generative Object detection, semantic segmentation, and change detection
SkySense [26] 2024 3 modalities HSROIs, TMsI, TSARI Swin, Vit Contrastive Scene classification, object detection, semantic segmentation, and change detection
OFA-Net [87] 2024 4 modalities 5 multi-modalities datasets ViT Generative Scene Classification and semantic segmentation
SAR-JEPA [38] 2024 SAR MSAR, SAR-Ship, SARSim, SAMPLE ViT Generative Target classification
MSFA [43] 2024 SAR ImageNet, DOTA, SARDet-100k CNN, ViT, Swin Generative Object detection
SARATR-X 2024 SAR ImageNet & 14 SAR datasets HiViT Generative Target classification and object detection
@@ 1. $\bullet$
    We propose a vision of general SAR target recognition, aiming to find universal solutions for various SAR ATR tasks and application in an era of SAR images with big unlabeled data and small labeled samples.
@@ 2. $\bullet$
    We have systematically investigated a foundation model solution for general SAR ATR and present the first SAR ATR foundation model named SARATR-X, which achieves new performance breakthroughs on a wide range of SAR target recognition datasets and settings.
@@ 3. $\bullet$
    We hope this promising work will stimulate research interest in SAR foundation models, thereby advancing SAR image interpretation with frontier AI technologies.
@@
The remainder of this paper is organized as follows. Sec. 2 introduces related work in remote sensing and SAR ATR. Sec. 3 introduces the proposed foundation model SARATR-X. Sec. 4 and 5 conducts extensive experiments to demonstrate the superiority of the proposed method. Sec. 6 concludes the paper and discusses future work.
2 Related Work
TABLE II: Related work about SSL and foundation models for SAR ATR. It can be found that the exploration of the foundation model in SAR ATR. Researchers have explored different dataset settings, CNN and Transfomer architectures, and various SSL algorithms. Inspired by their successful work, we formally present the first SAR ATR foundation model, SARATR-X, to achieve various recognition tasks.
Method Year Dataset Backbone SSL Tasks Description
RotANet [82] 2021 MSTAR CNN Generative Target classsification (MSTAR) Predicting rotational patterns of targets as a loss regularization term for classification task.
UACL [88] 2021 MSTAR, FUSAR-Ship CNN Contrastive Target classsification (MSTAR) Contrastive defense method to enhance model robustness to different adversarial attack.
PGIL [29] 2022 sea-ice, urban ResNet-18 Contrastive Scene classsification (sea-ice, urban) Contrastive learning between sub-frequency features of complex images and deep features of amplitude images.
BIDFC [92] 2022 MSTAR, OpenSARShip ResNet-18 Contrastive Target classsification (MSTAR, OpenSARShip) Weakly contrastive learning for pre-training in fine-grained datasets with similar sample.
TSCL [56] 2023 MSTAR ResNet-50 Contrastive Target classsification (MSTAR, OpenSARShip) Pre-processing of SAR images for noise removal to improve image quality for Contrastive learning.
FG-MAE [79] 2023 SSL4EO-S12 ViT-S Generative Scene classification (EuroSAT, BigEarthNet-MM), semantic segmentation (DFC2020) Hand-craft feature HOG as the target signal for MIM in SAR images.
SAR-JEPA [38] 2024 MSAR, SAR-Ship, SARSim, SAMPLE ViT-B Generative Target classification (MSTAR, FUSARShip, SAR-ACD) Local reconstruction and multi-scale gradient feature for MIM with small object in noise SAR images.
MSFA [43] 2024 ImageNet, DOTA, SARDet-100k Swin-B Generative Object detection (SARDet-100K, SSDD, HRSID) Multi-stage pre-training strategy for MIM from RGB images to SAR images.
SARATR-X 2024 ImageNet, 14 SAR target datasets HiViT-B Generative Target classification and Object detection (5 datasets and 8 settings) A first foundation model for target recognition in SAR Images.

### v2

# SARatrX: Towards Building A Foundation Model for SAR Target Recognition

@@ Abstract
Despite the remarkable progress in synthetic aperture radar automatic target recognition (SAR ATR), recent efforts have concentrated on the detection or classification of a specific and coarse category, e.g., vehicles, ships, airplanes, or buildings. One of the fundamental limitations of the top-performing SAR ATR methods is that the learning paradigm is supervised, task-specific, limited-category, closed-world learning, which depends on massive amounts of accurately annotated samples that are expensively labeled by expert SAR analysts and has limited generalization capability and scalability. In this work, we make the first attempt towards building a foundation model for SAR ATR, termed SARatrX. SARatrX learns generalizable representations via self-supervised learning (SSL) and provides a basis for label-efficient model adaptation to generic SAR target detection and classification tasks. Specifically, SARatrX is trained on 0.18 M unlabelled SAR target samples, which are curated by combining contemporary benchmarks and constitute the largest publicly available dataset till now. Considering the characteristics of SAR images, a backbone tailored for SAR ATR is carefully designed, and a two-step SSL method endowed with multi-scale gradient features was applied to ensure the feature diversity and model scalability of SARatrX. The capabilities of SARatrX are evaluated on classification under few-shot and robustness settings and detection across various categories and scenes, and impressive performance is achieved, often competitive with or even superior to prior fully supervised, semi-supervised, or self-supervised algorithms. Our SARatrX and the curated dataset are released at https://github.com/waterdisappear/SARatrX to foster research into foundation models for SAR ATR and SAR image interpretation.
@@
Synthetic Aperture Radar (SAR), Object Recognition, Object Detection, Foundation Model, Self-supervised learning, Deep Learning, Masked Image Modeling
https://github.com/waterdisappear/SARatrX
I Introduction
Synthetic aperture radar (SAR) [1, 2, 3, 4], developed from electromagnetic scattering in microwave frequency bands, plays a crucial role in active Earth observations, functioning effectively across diverse weather conditions and lighting environments. With the rapid development of SAR imaging techniques, high-resolution SAR images can be accessed more easily than before, enabling even more research opportunities for SAR image intelligent interpretation. SAR automatic target recognition (ATR), aiming at automatically localizing and classifying objects of interest (e.g., vehicles, ships, airplanes, or buildings) in SAR images, is a longstanding, important yet challenging problem in SAR image intelligent interpretation [5, 6, 7, 8]. SAR ATR plays an essential role in civil and national defense applications such as modern airport management, disaster management, urban planning and infrastructure monitoring, military reconnaissance, and maritime surveillance. Therefore, it has become an active research area for several decades [9, 10, 11, 12, 13, 14]. In the past decade, deep learning has brought tremendous success for SAR ATR [15, 5, 16, 17, 8]. Despite the significant progress, there are at least the following fundamental challenges that will need to be addressed to further advance the field of SAR ATR.
@@ I Introduction
Figure 1: The motivation and performance of SARatrX. (a) SAR ATR includes various imaging conditions, targets, scenes, and sensors. However, due to high costs, the datasets are often collected in specific settings for certain tasks. For example, MSTAR [18] is a ten-type vehicle target classification dataset in the X-band and grass scenarios, while SAR-Aircraft is a seven-type aircraft target detection dataset collected from three airports and a C-band satellite. Several specialized algorithms have been proposed for application to these datasets. However, the differing target characteristics, scene background information, and sensor parameters have complicated the generalization of existing algorithms. As such, this paper aims to develop a SAR ATR foundation model, a generalized method for conducting various tasks. (b) SARatrX performed well across 5 datasets with 8 settings. It was superior to existing SSL methods (BIDFC [19]) for target classification in the fine-grained vehicle MSTAR dataset [18] with a few-shot setting. In addition, it performed well under extended operating conditions (EOCs) [20] (i.e., imaging conditions with variable depression angle (EOCs-Depression), target configuration (EOCs-Config), and version (EOCs-Version)). SARatrX also demonstrated competitive object detection performance with existing supervised methods applied to various categories (SARDet-100K [21] and OGSOD [22]), as well as specific categories for ships (SSDD [23]) and aircraft (SAR-AIRcraft [24]). Our study shows the potential of a foundation model for generalized SAR ATR.
Firstly, task-specific property. One of the fundamental limitations of the current SAR ATR methods [25, 26, 27, 28, 29, 30] is that one model is trained and evaluated on one specific task. The detection and classification of a specific coarse category (e.g., vehicles, ships, airplanes, or buildings in Fig. 1.) all require their own deep models. As a result, the task-specific properties of these deep models pose significant challenges for training new tasks or developing a comprehensive SAR ATR system since each task must be learned independently from the ground up, requiring vast amounts of labeled data. This results in computational inefficiency, lower accuracy, and inconsistent results between the different models. Secondly, heavy reliance on supervised learning. Recent progress in SAR ATR, while substantial, has been limited to supervised learning, which heavily depends on massive amounts of accurately annotated target samples that are expensively labeled by expert SAR analysts and have limited generalization capability and scalability. However, the scarcity of expert SAR analysts cannot meet such an exhaustive requirement, leaving vast amounts of SAR images unlabelled and unexploited. Thirdly, the ignorance of SAR image characteristics in model designs. The imaging characteristics of SAR imagery differ significantly from those of optical imagery, leading to a significant domain gap between natural and SAR images. This raises significant challenges when one intends to transfer prior knowledge from the natural image domain. Different strong prior knowledge of SAR imagery, including speckle noise, discrete target appearances, and the lack of geometry, texture, and contour cues, needs special consideration when designing backbone architectures. Most of the current mainstream backbones and methods designed on natural images are not suitable for the aforementioned information. Finally, underdeveloped open-source ecosystem. Due to data sensitivity, the open-source ecosystem across the entire field is underdeveloped, making it challenging to share code and data publicly. Currently, there are no large and representative benchmark datasets for SAR ATR. As a result, this locks the potential of recent deep learning techniques for SAR ATR and significantly slows down the development of this field.
Recently, the remarkable success of foundation models (FMs) [31, 32, 33, 34] has led to a learning paradigm shift in artificial intelligence. Foundation models [35], pretrained on extensive data in a task-agnostic manner (generally via Self-supervised learning), can be flexibly adapted to a wide range of downstream tasks. Self-supervised learning (SSL) [36, 37, 7, 38] can be used to mitigate label inefficiencies by exploring supervision in the data directly, thereby reducing the reliance on expensive expert labeling while efficiently scaling the data and models. FMs shine in a broad range of areas, including natural language processing, computer vision, speech recognition, and medical image analysis. FMs have also been explored in remote sensing image understanding, but it is mostly limited to the evaluation of optical data. To our knowledge, the huge potential of FMs for SAR image interpretation remains completely locked.
In our preliminary work, a novel SSL method for SAR imagery named SAR Joint-Embedding Predictive Architecture (SAR-JEPA) [39] was proposed and demonstrated promising results. Given the aforementioned discussion, to fully unlock the potential of LMs for SAR image interpretation, we present the first attempt toward building a foundation model for SAR ATR and propose SARatrX that learns generalizable representations via SSL and provides a basis for label-efficient model adaptation to generic SAR target detection and classification tasks. While conceptually simple, existing methods for building FMs in other domains cannot be just directly applied to SAR ATR. We are still facing significant challenges, which we will discuss below and address with SARatrX.
@@
Pre-training datasets must include diverse target categories and imaging conditions to accommodate various downstream tasks. However, SAR ATR lacks a large-scale dataset, such as ImageNet [40], and the most common MSTAR dataset only includes fine-grained vehicle categories that are not suitable for larger-scale pre-training. As such, several SSL methods combine different datasets. Specifically, SARDet-100K [21] incorporates 9 SAR target detection sets. With the number of open-source datasets used for SAR ATR increasing, most of the open-source datasets were integrated into this study as part of the pre-training. A total of 14 classification and detection datasets with different target categories and imaging conditions were included in this study as a new pre-training dataset, SARDet-180K, to explore the potential of FMs, as seen in Table I.
@@
Model backbones aim to achieve better spatial representations in remote sensing images, especially for small target signatures in large imagery. Transformers and convolutional neural networks are among the most common architectures used for these tasks. Transformers offer better spatial resolution without downsampling, while convolutional neural networks (CNNs) provide higher efficiency through convolution kernels. The performance of different architectures was tested in MSFA [21] with SAR detection tasks, the results of which have shown the Swin transformer backbone to be the most effective for remote sensing images. As such, HiViT [41] was selected for use in this study, as it has the advantages of a Swin transformer and can drop patches in Masked Image Modeling (MIM).
@@
Self-supervised learning is complicated by SAR image quality, which is negatively affected by speckle noise in coherent imaging. The coherent imaging resulting visual features are also not as distinct or rich as in natural RGB images. Contrastive learning [19, 42] uses data augmentation and preprocessing to reduce noise, while MIM [43, 39, 21] applies various target feature for guided signals to suppress noise. As such, the primary task of SAR SSL is the enhance the quality of feature learning and guide signals. For example, PGIL [6] leveraged a sub-frequency feature of complex SAR images to learn physics information, while our SAR-JEPA [39] applied multi-scale gradient ratios to solve for multiplicative speckle noise and capture target shapes. Furthermore, multi-stage training [21] from ImageNet to SAR diminished the effects of noise on model diversity, as seen in Fig. 4. Thus, we applied two-step pretraining from ImageNet to SAR to increase model diversity during pretraining with SAR images. Besides, multiscale gradient features were used as high-quality guide signals for MIM with SAR images.
@@
Evaluation tasks need to comprehensively evaluate the performance of a foundation model for different tasks and settings. The availability of 3 open-source target datasets was utilized by first constructing a fine-grained classification dataset, SAR-VSA, with 25 categories to evaluate the effectiveness of the proposed improvements. A comprehensive comparison was then performed between the proposed SARatrX and existing methods for public classification and detection tasks.
SARatrX achieved superior performance in 5 datasets across 8 task settings, as shown in Fig. 1, which is competitive with prior methods on various SAR ATR tasks (few-shot vehicle classification, robust classification, detection with specific categories or various categories). We hope that this work could advance the development of the intersection of general SAR target recognition and FMs. The primary contributions of this study can be summarized as follows:
@@ 1. $\bullet$
    We present the first foundation model called SARatrX, which learns generalizable representations via SSL from large-scale unlabelled data and provides a basis for generic SAR target detection and classification tasks.
@@ 2. $\bullet$
    We systematically investigate a foundation model framework for SAR ATR. We build a currently the largest publicly available pre-training dataset SARdet-180k and fully discussed the model architecture and proposed SSL method with many comparisons.
@@ 3. $\bullet$
    The ability of SARatrX is comprehensively evaluated with various SAR ATR tasks, such as few-shot classification, robust classification, ship detection, aircraft detection, and detection with various categories tasks.
@@
The remainder of this paper is organized as follows. Sec. II introduces related work in remote sensing and SAR ATR. Sec. III discusses the proposed foundation model (SARatrX). Secs. IV and V conduct extensive experiments to demonstrate the superiority of the proposed method. Sec. VI concludes the paper and discusses future work.
II Related Work